Find the section of the conicoid (x2 /2) − (y 2/ 3 )= 2z by the plane x− 2y+ z = 1. What conic does this section represent? Justify your answer.


1
Expert's answer
2021-06-11T16:38:31-0400

Let us find the section of the conicoid "\frac{x^2}{2} −\frac{y^2}{3}= 2z"  by the plane x− 2y+ z = 1. It follows that "\frac{x^2}{2} −\frac{y^2}{3}= 2(1-x+2y)," and hence "\frac{x^2}{2} −\frac{y^2}{3}+2x-4y=2." It is equivalent to "3x^2-2y^2+12x-24y=12," and hence "3(x^2+4x)-2(y^2+12y)=12." We conclude that "3(x+2)^2-2(y+6)^2=-48," and hence the equation of the section is "\frac{(y+6)^2}{24}-\frac{(x+2)^2}{16}=1."

It follows that this section represents a hyperbola.